OpenStudy (idealist10):

Simplify: (1/6)ln abs(y+1)-(1/2)ln abs(y-1)+(1/3)ln abs(y-2).

OpenStudy (idealist10):

Please help me.

OpenStudy (anonymous):

Is this: \[\frac{1}{6}\ln (|y+1|) - \frac{1}{2} \ln(|y-1|) + \frac{1}{3}\ln (|y-2|)\]

OpenStudy (idealist10):

Yes.

OpenStudy (anonymous):

Actually I don't know how to reach there but let me provide what I know: \[a \ln(x) = \ln(x^a)\] \[\ln(x) + \ln(y) = \ln(xy)\] \[\ln(x) - \ln(y) = \ln(\frac{x}{y})\]

OpenStudy (anonymous):

You should try to use these formulae there..

OpenStudy (idealist10):

So I got (y+1)^(1/6)*(y-1)^(-1/2)*(y-2)^(1/3), is this right?

OpenStudy (anonymous):

But this seems to be going nowhere.. :( Sorry..

OpenStudy (idealist10):

How can I simplify this more?

OpenStudy (idealist10):

@myininaya

OpenStudy (anonymous):

\[\frac{1}{6}\ln|y+1| - \frac{1}{2} \ln|y-1| + \frac{1}{3}\ln |y-2|\] Factor out \(\dfrac{1}{6}\) and apply the properties: \[\frac{1}{6}\left(\ln|y+1| - \frac{6}{2} \ln|y-1| + \frac{6}{3}\ln |y-2|\right)\\ \frac{1}{6}\left(\ln|y+1| - \ln|y-1|^3 + \ln (y-2)^2\right)\\ \frac{1}{6}\ln\left|\frac{(y+1)(y-2)^2}{(y-1)^3}\right|\]